It costs Brandon $430 to maintain his car this year. Brandon's mechanic says that he can expect his maintenance cost to increase by 15% per year for the next 8 years.
A. Write a function for the maintenance cost for Brandon's car after n years.
B. When will Brandon need to spend $1000 to maintain his car for the year?
C. Use your function to predict how much it will cost Brandon to maintain his car in 8 years.

Answer by stanbon(75887)

It costs Brandon $430 to maintain his car this year. Brandon's mechanic says that he can expect his maintenance cost to increase by 15% per year for the next 8 years.
A. Write a function for the maintenance cost for Brandon's car after n years.
M(n) = 430*1.15^n
-----------------------------
B. When will Brandon need to spend $1000 to maintain his car for the year?
Solve:
430*1.15^n = 1000
1.15^n = 2.3256
n*log(1.15) = log(2.3256)
n = log(2.3256)/log(1.15) = 6.04 years
-------------------------------------------
C. Use your function to predict how much it will cost Brandon to maintain his car in 8
M(8) = 430*1.15^8 = $1315.38
==================
Cheers,
Stan H.
==================